OpenStudy (anonymous):

I normally don't have trouble with homework assignments, but this one i have no idea how to solve. Rectangle R has varying length l and width w but with a constant perimeter of 4 feet a) Express the area A as a function of length l. Write a list of everything you know about this function. b) For what values of l and w will the area of R be the greatest? Give an algebraic explanation. Then give a geometric explanation. Be sure to include a graph with relevant points labeled. Can any one help me find out how to solve this problem?

OpenStudy (phi):

start by writing an equation for finding the perimeter. can you do that ?

OpenStudy (anonymous):

Heres my work so far W+L=2 L=2-W And i can plug this into A=L*W A=(2-W)*W A=2W*W^2 But thats where i get stuck...Am i doing something wrong? Thank you for helping me

OpenStudy (phi):

the last line has a typo? * should be -

OpenStudy (anonymous):

oops, yes, thats supposed to be -. But im still stuck on what to do next

OpenStudy (phi):

if we use x for w and y for A, we can write it as y= -x^2 +2x that is the equation of a parabola (see attached)

OpenStudy (phi):

As you can see, the vertex (where it peaks) occurs when x =1 i.e. when the width is 1 btw, for part (a) they want a function of L, and you found a function of W you should re-do that part to get f(L)= L(2-L) = -L^2 +2L (same parabola, so we get the same peak)

OpenStudy (anonymous):

ohhhhh, Thats how to graph it. Thank you, So to see what values would be the greatest i would use the graph? (1,1)????

OpenStudy (phi):

I expect they want you to find the vertex of the parabola one way is to change to vertex form y = a(x-h)^2 +k where (h,k) is the vertex you do that by completing the square.

OpenStudy (phi):

the other way is to remember that when in standard form y = a x^2 + bx + c the x coordinate of the vertex is at x= -b/(2a)

OpenStudy (phi):

graphing is one way to find the vertex, but to graph the parabola, people normally find the vertex using one of the ideas up above. (unless you use a computer to graph it) to help make the graph.

OpenStudy (anonymous):

Thank you for helping me with this! If its not to much trouble, can you help me with the second part? This time rectangle R has varying length l and width w but with a constant area of 4 square feet. a) Express the perimeter P as a function of length l. What type of function is P? What is the domain of P? b) Describe the asymptotic behavior of P. What can you say about rectangle R because of this behavior? Could you have made a similar statement about R back in Task 1? c) For what values of l and w will the perimeter of R be theleast? Give a geometric explanation. Be sure to include a graph with relevant points labeled.

OpenStudy (phi):

how far did you get ?

OpenStudy (anonymous):

I haven't started yet, I never got past part one. Im attempting to solve it now.

OpenStudy (anonymous):

I think i just have to do the same thing, just for the perimeter this time. Am I right?

OpenStudy (phi):

yes, it is the same approach.

OpenStudy (anonymous):

P=2(L+W) Area=L*W=4ft^2 W=4/L P=2(L+4/L) Is that right?

OpenStudy (phi):

yes. if we relabel L as x and P as y this is the same as y = 2x +8/x that is a hyperbola, but I'm not sure how to prove it. If you download Geogebra http://www.geogebra.org/cms/en/ you can make this plot

OpenStudy (anonymous):

Ok Thanks for your help!

OpenStudy (phi):

yw